### 08 Mar 2026 v3

# Tripartite information of free fermions: a universal entanglement coefficient from the sine kernel

Abstract
We study the tripartite information $I_{3}$ of free fermions on two-dimensional lattices partitioned into three adjacent strips of width $w$. Translation invariance yields the exact decomposition $I_{3}=\sum_{k_{y}}g(k_{F}(k_{y})\,w)$, where $g(z)$ is a universal function of the scaling variable $z=k_{F}w$, determined by the spectrum of the sine-kernel (Slepian) integral operator. We find that $g(z)$ has a unique zero at $z^{*}=1.329\pm 0.001$: modes with $k_{F}w<z^{*}$ violate monogamy of mutual information ($g>0$), while modes with $k_{F}w>z^{*}$ satisfy it ($g<0$).
The central analytical result is $g(z)=cz+O(z^{3}\ln z)$ with $c=3\ln(4/3)/\pi\approx 0.2747$, derived from the rank-1 limit of the sine kernel. Two exact cancellations—of the $z\ln z$ area-law terms and of the $z^{2}$ terms—are intrinsic to the $I_{3}$ combination. The coefficient $c$ generalizes to $n$-partite information: $c_{n}=(n/\pi)\ln R_{n}$ with $R_{n}$ a rational number from binomial combinatorics. For Rényi entropy of index $\alpha$, we show that $g_{\alpha}(z)\sim z^{\alpha}$ for $\alpha<2$ and $g_{2}(z)=-(8/\pi^{3})z^{3}$: von Neumann entropy ($\alpha=1$) uniquely gives linear sensitivity to Lifshitz transitions, while Rényi-2 gives only cubic sensitivity. We verify all predictions on square, triangular, and cubic lattices. DMRG calculations on the interacting $t$-$V$ model provide evidence that the linear coefficient is independent of the Luttinger parameter $K$.
1 Introduction
The tripartite information $I_{3}(A{:}B{:}D)=S_{A}+S_{B}+S_{D}-S_{AB}-S_{AD}-S_{BD}+S_{ABD}$ quantifies the structure of multipartite correlations in quantum states [1]. When $I_{3}\leq 0$ for all subsystems, the state satisfies monogamy of mutual information (MMI)—a property guaranteed by holographic duality [1, 2] but violated by free field theories [3, 4].
For lattice fermion systems, the status of $I_{3}$ has remained unclear. Calabrese, Cardy, and Tonni computed $I_{3}$ for disjoint intervals in 1+11{+}1D CFTs [5, 6]. Parez et al. [7] obtained multipartite information for free fermions on Hamming graphs. Agón, Bueno, and Casini [4] analyzed the long-distance behavior of $I_{3}$ in generic CFTs. However, a systematic study of $I_{3}$ as a function of Fermi surface geometry—and in particular, the conditions under which MMI holds or fails—has been lacking.
In this work we present a complete analytical framework for $I_{3}$ of free fermions on two-dimensional lattices, partitioned into three adjacent strips of width $w$. The key results are:
(i) An exact decomposition $I_{3}=\sum_{k_{y}}g(k_{F}(k_{y})\,w)$ over transverse modes, with a universal function $g(z)$ determined by the sine-kernel spectrum. The function has a unique zero at $z^{*}=1.329$, separating MMI-violating modes ($z<z^{*}$) from MMI-satisfying modes ($z>z^{*}$). Consequently, MMI is scale-dependent: any metallic state satisfies it at large $w$ and violates it at small $w$.
(ii) An analytical derivation of $g(z)=cz+O(z^{3}\ln z)$ with $c=3\ln(4/3)/\pi$, from the rank-1 structure of the sine kernel at small $z$. The result rests on two exact cancellations ($\sum a_{n}n=\sum a_{n}n^{2}=0$) intrinsic to the $I_{3}$ combination.
(iii) A generalization to $n$-partite information ($c_{n}=(n/\pi)\ln R_{n}$) and to Rényi entropy, showing that von Neumann entropy is the unique index giving linear sensitivity to Lifshitz transitions.
(iv) Verification on square, triangular, and cubic lattices, including the $t^{\prime}$-driven sign change, width dependence, and directional anisotropy.
The paper is organized as follows. Section 2 presents the $k_{y}$-decomposition and properties of $g(z)$. Section 3 derives the linear coefficient $c$. Sections 4 and 5 generalize to $n$-partite and Rényi cases. Section 6 derives physical consequences for Lifshitz transitions. Section 7 presents numerical verification. Section 8 discusses implications and open questions.
2 Framework: mode decomposition and universal function
2.1 Model
Consider free fermions on an $L\times L$ lattice with periodic boundary conditions and dispersion $\varepsilon(\bm{k})$. On the square lattice: $\varepsilon(\bm{k})=-2t_{x}\cos k_{x}-2t_{y}\cos k_{y}-4t^{\prime}\cos k_{x}\cos k_{y}$; on the triangular lattice: $\varepsilon(\bm{k})=-2t(\cos k_{x}+\cos k_{y}+\cos(k_{x}-k_{y}))$. The system is partitioned into three adjacent strips $A$, $B$, $D$ of width $w$ in the $x$-direction.
All entropies are computed from the single-particle correlation matrix via the Peschel formula [8, 18]: $S_{X}=-\mathrm{Tr}[C^{(X)}\ln C^{(X)}+(1-C^{(X)})\ln(1-C^{(X)})]$. Translation invariance in $y$ block-diagonalizes over $k_{y}$ modes.
2.2 $k_{y}$-decomposition
Figure 1: Fermi surfaces at half filling for: (a) square lattice, $t^{\prime}=0$ (perfect nesting); (b) $t^{\prime}=0.12\approx t^{\prime}_{*}$; (c) $t^{\prime}=0.30$; (d) triangular lattice. The value of $I_{3}/L$ ($L=256$, $w=2$) is shown below each panel. The sign change between (a) and (b) is controlled by the distribution of $k_{F}(k_{y})$ relative to the universal constant $z^{*}/w$.
The tripartite information decomposes exactly:
$I_{3}=\sum_{k_{y}}I_{3}^{(1D)}(k_{F}(k_{y}),\,w),$ (1)
where $I_{3}^{(1D)}(k_{F},w)$ is the tripartite information of three adjacent blocks of width $w$ in a one-dimensional chain with Fermi momentum $k_{F}$. In the thermodynamic limit, $I_{3}^{(1D)}$ is computed exactly from the Toeplitz matrix [9]:
$I_{3}^{(1D)}(k_{F},w)=3\,S(T_{w})-2\,S(T_{2w})-S(T_{w}^{(AD)})+S(T_{3w}),$ (2)
where $T_{\ell}$ is the $\ell\times\ell$ Toeplitz matrix with entries $C_{ij}=\sin(k_{F}|i{-}j|)/(\pi|i{-}j|)$ and $T_{w}^{(AD)}$ is the $2w\times 2w$ correlation matrix for the disjoint block $A\cup D$.
2.3 Universal function $g(z)$
In the scaling limit $w\to\infty$ at fixed $z=k_{F}w$, the Toeplitz matrices converge to the sine-kernel integral operator [10]:
$K_{z}(x,y)=\frac{\sin[z(x-y)]}{\pi(x-y)},$ (3)
and $I_{3}^{(1D)}\to g(z)$, a universal function of $z$ alone. For a two-dimensional system:
$I_{3}=\frac{L}{2\pi}\int g(k_{F}(k_{\perp})\,w)\,dk_{\perp}.$ (4)
The key properties of $g(z)$, established by analytical arguments at small and large $z$ and by systematic numerical evaluation at intermediate $z$ (see Appendix A), are:
Result. (a) $g(0)=0$ and $g^{\prime}(0^{+})=c>0$ (analytical). (b) $g$ has a unique maximum at $z_{\max}\approx 0.56$, with $g(z_{\max})\approx 0.088$ (numerical). (c) $g$ has a unique zero at $z^{*}=1.329\pm 0.001$ (numerical; see below for error estimate). (d) $g(z)<0$ for all $z>z^{*}$ (numerical for $z\in(z^{*},10)$; analytical for $z\to\infty$).
The zero $z^{*}$ separates two regimes. For $z<z^{*}$, the direct mutual information I(A:D)I(A{:}D) across the gap $B$ exceeds the concavity deficit $\Delta^{2}S$ of single-interval entropy, giving $g>0$. For $z>z^{*}$, entropy concavity dominates, giving $g<0$. The zero is determined to $0.12\%$ by the cancellation of only two Slepian eigenvalue contributions, with exponentially decaying tail contributions bounded by prolate spheroidal eigenvalue estimates [10].
Figure 2: (a) Analytical $I_{3}^{(1D)}(\nu)$ from Eq. (2) for strip widths w=1w=11616. (b) The product $k_{F}^{*}w$ converges to $z^{*}\approx 1.329$ (dashed red). (c) Scaling collapse: $I_{3}$ vs. $z=k_{F}w$ for various $w$, confirming convergence to a universal curve $g(z)$.
2.4 Scale-dependent MMI
Since $g(z)<0$ for $z>z^{*}$, any mode with $k_{F}w>z^{*}$ contributes negatively to $I_{3}$. As $w$ increases, eventually all modes satisfy this condition, and $I_{3}<0$. Conversely, modes with small $k_{F}$ (near a Fermi surface touching point or at low filling) have $z<z^{*}$ and contribute $g>0$. Thus:
MMI is not a property of the quantum state alone, but of the pair (state, observation scale). Any metallic state violates MMI for sufficiently narrow strips and satisfies it for sufficiently wide ones, provided the Fermi surface contains modes with small $k_{F}$.
3 Derivation of the linear coefficient
3.1 Rank-1 structure at small $z$
At small $z=k_{F}w$, the sine kernel becomes approximately constant: $\mathrm{sinc}(k_{F}d)=1+O(z^{2}/w^{2})$ for all site separations $d$ within the $3w$-site region. The correlation matrix reduces to rank 1, with a single nonzero eigenvalue $\lambda_{0}^{(n)}=nz/\pi$ for a block of $nw$ sites. This follows from the Slepian eigenvalue asymptotics $\lambda_{m}\sim(z/\pi)^{2m+1}$ for $z\to 0$ [10], which ensures exponential suppression of all $m\geq 1$ modes at small $z$.
Crucially, the disjoint block $A\cup D$ has the same leading eigenvalue $\lambda_{0}^{(AD)}=2z/\pi+O(z^{3})$, since a constant kernel cannot distinguish contiguous from separated sites. The entropy of each block is therefore $S_{n}=h(nz/\pi)+O(z^{3}\ln z)$ and $S_{AD}=h(2z/\pi)+O(z^{3}\ln z)$, where $h(\lambda)=-\lambda\ln\lambda-(1{-}\lambda)\ln(1{-}\lambda)$ is the binary entropy.
3.2 Two exact cancellations
With these expressions, the four-term structure of $g(z)$ collapses to three terms:
$g(z)=\sum_{k=1}^{3}a_{k}\,h(kz/\pi)+O(z^{3}\ln z),$ (5)
with inclusion-exclusion coefficients $a_{k}=(3,-3,1)$ for $k=(1,2,3)$. Expanding $h(\lambda)=-\lambda\ln\lambda+\lambda-\lambda^{2}/2+O(\lambda^{3})$:
$g(z)=-\frac{z}{\pi}\sum_{k}a_{k}k\ln\frac{kz}{\pi}+\frac{z}{\pi}\sum_{k}a_{k}k-\frac{z^{2}}{2\pi^{2}}\sum_{k}a_{k}k^{2}+O(z^{3}\ln z).$ (6)
First cancellation: $\sum a_{k}k=3-6+3=0$. This eliminates both the explicit constant and the $\ln(z/\pi)$ piece from the logarithmic sum, leaving only $\sum a_{k}k\ln k$.
Second cancellation: $\sum a_{k}k^{2}=3-12+9=0$. This eliminates the $z^{2}$ term from the $\lambda^{2}/2$ correction in $h(\lambda)$.
What survives is a single combination of logarithms:
$\sum_{k}a_{k}k\ln k=\ln\prod_{k}k^{a_{k}k}=\ln\frac{1^{3}\cdot 3^{3}}{2^{6}}=-3\ln\frac{4}{3}.$ (7)
Therefore:
$\boxed{g(z)=cz+O(z^{3}\ln z),\qquad c=\frac{3\ln(4/3)}{\pi}=0.274716\ldots}$ (8)
3.3 Higher-order structure
The first correction to linearity is not $z^{2}$ (killed by the second sum rule) but $z^{3}\ln z$, from the second Slepian eigenvalue $\lambda_{1}\sim 0.274\,(z/\pi)^{3}$ [10]:
$g(z)=cz+Dz^{3}\ln z+Ez^{3}+O(z^{5}\ln z),$ (9)
with $D\approx 0.218$ and $E\approx-0.226$ (determined numerically), and $\sum a_{k}k^{3}=3-24+27=6\neq 0$.
Figure 3: The ratio $g(z)/z$ on a log-linear scale (main panel) and $g(z)$ itself (inset), computed at $w=64$. At small $z$, the ratio converges to $c=3\ln(4/3)/\pi\approx 0.2747$ (dashed red), confirming the rank-1 derivation. The zero at $z^{*}\approx 1.329$ (dotted green) separates the pocket regime ($g>0$) from the monogamy regime ($g<0$).
This hierarchy has physical content. The leading Slepian eigenmode encodes only the Fermi surface area (via Cavalieri’s principle): the two identities $\sum a_{k}k=\sum a_{k}k^{2}=0$ eliminate all nonlinear information from this mode. Shape information enters only through the second eigenmode, producing the nonanalytic $O(z^{3}\ln z)$ correction.
4 $n$-partite generalization
The structure generalizes to $n$-partite information $I_{n}$ of $n$ adjacent strips. The inclusion-exclusion coefficients are $a_{k}=(-1)^{k+1}\binom{n}{k}$, and the rank-1 analysis gives:
$c_{n}=\frac{n}{\pi}\ln R_{n},\qquad R_{n}=\prod_{j=0}^{n-1}(j{+}1)^{(-1)^{j+1}\binom{n-1}{j}}$. (10)
The first values are:
$c_{2}=\frac{\ln 4}{\pi},\quad c_{3}=\frac{3\ln\frac{4}{3}}{\pi},\quad c_{4}=\frac{4\ln\frac{32}{27}}{\pi},\quad c_{5}=\frac{5\ln\frac{3125}{2592}}{\pi}.$ (11)
The mutual information case $c_{2}=\ln 4/\pi\approx 0.441$ [15] is the first member of this family.
The sum rules extend: $\sum_{k}a_{k}k=0$ for all $n\geq 2$ (eliminating $z\ln z$), and $\sum_{k}a_{k}k^{2}=0$ for all $n\geq 3$ (eliminating $z^{2}$). For $n=2$, $\sum a_{k}k^{2}=2\neq 0$, so $I_{2}$ has a $z^{2}$ correction that $I_{3}$ and all higher multipartite measures lack. This is a structural advantage of the tripartite combination: it is the simplest multipartite measure with both cancellations.
5 Rényi uniqueness of von Neumann entropy
5.1 Analytical argument
For Rényi-$\alpha$ entropy, the single-mode entropy at small $\lambda$ is:
$h_{\alpha}(\lambda)=\frac{\alpha}{\alpha-1}\lambda-\frac{\alpha}{2}\lambda^{2}+O(\lambda^{3})\qquad(\alpha>1).$ (12)
The leading term is analytic, so $\sum a_{k}k=0$ kills it completely. No $z\ln z$ term is generated, and $c_{\alpha}\equiv\lim_{z\to 0}g_{\alpha}(z)/z=0$ for all $\alpha>1$.
For $\alpha=2$, the first surviving term comes from $O(\lambda^{3})$ in $h_{2}$, giving:
$g_{2}(z)=-\frac{8}{\pi^{3}}\,z^{3}+O(z^{5}),\qquad\text{where }-\frac{8}{\pi^{3}}=\frac{-4}{3\pi^{3}}\sum a_{k}k^{3}=\frac{-4\cdot 6}{3\pi^{3}}.$ (13)
For non-integer $\alpha$ with $1<\alpha<2$, the entropy $h_{\alpha}(\lambda)$ has a nonanalytic $\lambda^{\alpha}$ contribution that is not killed by the sum rule, giving $g_{\alpha}(z)\sim z^{\alpha}$. For $\alpha<1$, the $\lambda^{\alpha}$ term dominates and $g_{\alpha}(z)/z$ diverges.
Only for $\alpha=1$ (von Neumann) does the $\lambda\ln\lambda$ nonanalyticity survive the first sum rule to produce a finite, nonzero linear coefficient $c$.
5.2 Numerical verification
Table 1: Rényi scaling of $g_{\alpha}(z)\sim z^{\beta}$. Exponents extracted by linear fit of $\ln|g_{\alpha}(z)|$ vs. $\ln z$ over $z\in[10^{-3},10^{-2}]$ using Toeplitz matrices at $w=80$. Fit residuals are below $1\%$ for all $\alpha$. The $\alpha=2$ coefficient $8/\pi^{3}\approx-0.258$ is confirmed to $0.15\%$.
$\alpha$ $\beta$ (num.) $\beta$ (pred.) Coefficient Note
$0.5$ $0.49$ $1/2$ +0.52+0.52 superlinear
11 $1.00$ 11 +0.2747+0.2747 $3\ln(4/3)/\pi$
$1.5$ $1.48$ $3/2$ +0.09+0.09 subanalytic
22 $3.02$ 33 0.258-0.258 $8/\pi^{3}$
33 $4.0$ >3>3 0.005-0.005 higher cancel.
5.3 The Rényi-$\alpha$ function at finite $z$
Figure 4: Rényi-index dependence of the scaling exponent $\beta$ in $g_{\alpha}(z)\sim z^{\beta}$. Only $\alpha=1$ (von Neumann, star) gives $\beta=1$ and hence linear sensitivity to Lifshitz transitions. For $\alpha>1$, the analytical linear term is killed by $\sum a_{k}k=0$, and only the nonanalytic piece survives. Dashed line: $\beta=\alpha$ for $\alpha<2$.
For $\alpha>1$, the function $g_{\alpha}(z)$ has qualitatively different behavior from the von Neumann case: it oscillates with multiple sign changes. For $\alpha=2$ (experimentally relevant [16, 17]), $g_{2}(z)$ is negative for $z<1.19$, positive for $1.19<z<2.11$, and oscillates thereafter. This is a prediction testable in cold-atom experiments.
5.4 Physical consequence
The pocket signal at a Lifshitz transition scales as $\Delta I_{3}\propto g^{\prime}(0^{+})\cdot\delta$. For von Neumann: $\Delta I_{3}^{(\alpha=1)}\propto c\,\delta$. For Rényi-2: $\Delta I_{3}^{(\alpha=2)}\propto\delta^{3}$—suppressed by a factor $\delta^{2}$ relative to the von Neumann signal. This means experimental probes measuring $S_{2}$ [16, 17] would require correspondingly higher precision to detect the leading Lifshitz singularity.
6 Physical consequences
The linear coefficient $c$ implies exact predictions for entanglement singularities at Lifshitz transitions [11].
Pocket transition. In 2D, when a small electron pocket appears at $\mu>\mu_{c}$, all pocket modes have $k_{F}w\ll z^{*}$ and contribute $g(k_{F}w)\approx c\,k_{F}w$. Evaluating the integral over the elliptical pocket:
$\Delta I_{3}=\frac{cwm}{2}\,L\,\delta\,\Theta(\delta),\qquad\delta=\mu-\mu_{c},$ (14)
with $m=\sqrt{m_{x}m_{y}}$ the geometric-mean effective mass. The derivative $\partial I_{3}/\partial\mu$ has a discontinuity at $\mu_{c}$, mirroring the 2D density-of-states discontinuity with the entanglement coefficient $c$ replacing the DOS.
Neck transition (VHS). At a van Hove singularity, modes near the saddle have $k_{F}(q)=\gamma\sqrt{q^{2}+\delta/\beta}$, and for sufficiently small $w$ (so that all neck modes have $k_{F}w\ll z^{*}$), $g^{\prime}(k_{F}w)\approx c$:
$\frac{\partial I_{3}}{\partial\mu}=-\frac{cLw\gamma}{4\pi\beta}\ln|\mu-\mu_{c}|+\mathrm{const},$ (15)
a logarithmic divergence with fully determined coefficient.
Entanglement tomography. The angular dependence $I_{3}(\theta)$ for strips at angle $\theta$ probes Fermi surface shape. For an $n$-fold deformation $k_{F}(\phi)=\bar{k}_{F}(1+\varepsilon\cos n\phi)$, the Widom coefficient responds at $O(\varepsilon^{2})$ (perimeter is quadratic in $\varepsilon$ at fixed area). In contrast, $I_{3}(\theta)$ is expected to respond at $O(\varepsilon)$ through the nonlinearity of $g(z)$, since the integral $\int g(k_{F}(\phi)\,w)\,d\phi$ is generically linear in $\varepsilon$ when $g$ is nonlinear. A quantitative analysis of this sensitivity is left for future work.
7 Numerical verification
All free-fermion results are computed via the Peschel formula applied to $3w\times 3w$ Toeplitz matrices at w=64w=648080. Finite-$w$ corrections are below $0.1\%$ for all quoted values, as verified by comparing $w=64$ and $w=128$. Lattice calculations use $L=256$ unless noted otherwise.
7.1 The constant $c$
Table 2: Convergence of $g(z)/z$ to $c=3\ln(4/3)/\pi=0.274716$ at $w=80$.
$z$ $g(z)/z$ Deviation from $c$ Note
0.001 0.274714 $0.0006\%$ 6 sig. figs
0.010 0.274593 $0.04\%$ $O(z^{2})$ onset
0.100 0.267240 $2.7\%$ $O(z^{3}\ln z)$
7.2 Sign change vs. $t^{\prime}$ on the square lattice
On the square lattice at half filling, $I_{3}>0$ at $t^{\prime}=0$ (diamond FS with nested modes at $k_{F}\to 0$) and $I_{3}<0$ for $t^{\prime}>t^{\prime}_{*}\approx 0.10$ (FS deformed away from nesting).
Table 3: $I_{3}/L$ vs. $t^{\prime}$ for the square lattice at half filling, $w=2$.
$t^{\prime}$ $I_{3}/L$ $t^{\prime}$ $I_{3}/L$
0.00 +0.00606+0.00606 $I_{3}>0$ 0.12 0.00162-0.00162 $I_{3}<0$
0.05 +0.00407+0.00407 0.20 0.00786-0.00786
0.08 +0.00171+0.00171 0.30 0.01439-0.01439
0.10 $\approx 0$ crossing 0.50 0.02162-0.02162
The mechanism is precisely the $k_{y}$-decomposition: at $t^{\prime}=0$, the diamond FS gives $k_{F}(k_{y})=\pi-|k_{y}|$, which vanishes at $k_{y}=\pm\pi$. These modes with $k_{F}w<z^{*}$ contribute $g>0$ and dominate $I_{3}$. Increasing $t^{\prime}$ shifts all $k_{F}(k_{y})$ away from zero, entering the $g<0$ regime.
7.3 Square vs. triangular lattice
Table 4: $I_{3}/L$ for square vs. triangular lattice at half filling, $L=256$. The triangular lattice has $k_{F}(k_{y})$ bounded away from zero, so all modes have $z>z^{*}$ at moderate $w$.
Lattice $w=2$ $w=4$ $w=8$
Square +0.00606+0.00606 +0.00413+0.00413 +0.00228+0.00228
Triangular 0.02156-0.02156 0.00628-0.00628 0.00126-0.00126
On the triangular lattice, $I_{3}$ becomes positive near the van Hove singularity ($\nu\approx 0.71$), where the FS passes through a saddle point creating modes with $k_{F}\to 0$.
7.4 Universal zero-crossing constant
Table 5: Convergence of $k_{F}^{*}w\to z^{*}\approx 1.329$.
$w$ $\nu_{*}$ $k_{F}^{*}w$ $w$ $\nu_{*}$ $k_{F}^{*}w$
2 0.1935 1.2160 16 0.0264 1.3267
4 0.1032 1.2969 32 0.0132 1.3283
8 0.0525 1.3205 64 0.0066 1.3287
7.5 Width dependence and directional anisotropy
Table 6: Directional $I_{3}$ for anisotropic hopping, $L=256$, $w=2$.
$t_{x}$ $t_{y}$ $I_{3}^{(x)}/L$ $I_{3}^{(y)}/L$ $|I_{3}^{(x)}/I_{3}^{(y)}|$
1.0 1.0 +0.0061+0.0061 +0.0061+0.0061 1.0
1.5 0.5 0.0201-0.0201 0.0025-0.0025 8.1
2.0 0.5 0.0222-0.0222 0.0013-0.0013 17.2
7.6 Pocket formula
Table 7: Pocket formula $\Delta I_{3}/(Lw\delta)\to cm/2$ at small $z_{\max}=\sqrt{2m\delta}\,w$.
[…]
2 $10^{-6}$ 0.003 0.27470 0.27472
4 $10^{-6}$ 0.006 0.54935 0.54943
8 $^{-6} 0.011$ 1.09837 1.09886
8 Discussion
Origin of $c$. The constant $c=3\ln(4/3)/\pi$ arises from the interplay of three ingredients: the rank-1 limit of the prolate spheroidal operator, the inclusion-exclusion combinatorics of $I_{3}$, and the $\lambda\ln\lambda$ nonanalyticity of von Neumann entropy. The $n$-partite generalization $c_{n}=(n/\pi)\ln R_{n}$ reveals systematic combinatorial structure in the product representations $R_{n}$.
Von Neumann uniqueness. The uniqueness of $\alpha=1$ for linear Lifshitz sensitivity is unexpected and experimentally relevant. Current cold-atom experiments measure Rényi-2 entropy [16, 17], which gives only $O(\delta^{3})$ pocket sensitivity. Direct measurement of von Neumann $I_{3}$—or development of protocols to extract it from Rényi data—would enable linear detection of Fermi surface topology changes.
Widom coefficient vs. $I_{3}$. The Widom coefficient $c_{1}$ in $S_{A}\sim c_{1}L\ln L$ [12, 13, 14] depends only on the FS perimeter and is always positive. In contrast, $I_{3}$ changes sign as the FS geometry varies: Table 3 shows that $I_{3}$ reverses from positive to negative as $t^{\prime}$ increases, detecting the disappearance of nested modes with $k_{F}w<z^{*}$ that the Widom coefficient cannot distinguish. More generally, $I_{3}(\theta)$ for strips at angle $\theta$ may be sensitive to FS shape deformations at $O(\varepsilon)$, while the Widom coefficient responds only at $O(\varepsilon^{2})$. The $z^{2}$ cancellation ($\sum a_{k}k^{2}=0$) ensures that this sensitivity comes from the higher-order structure of $g(z)$ rather than trivial area scaling.
Interactions. The rank-1 derivation of Section 3 relies on the correlation matrix becoming approximately constant over the block at small $z$. For free fermions, this follows from $\mathrm{sinc}(k_{F}d)\approx 1$ when $k_{F}d\ll 1$. In a Luttinger liquid, the single-particle correlator acquires an anomalous exponent, $\langle c^{\dagger}_{i}c_{j}\rangle\sim\sin(k_{F}r)/r^{(K+K^{-1})/2}$, and the rank-1 approximation no longer holds in general. However, the rank-1 eigenvalue $\lambda_{0}=\sum_{j}C_{0j}/w$ is determined by the integral of the correlator over the block, which equals the mean occupation $k_{F}/\pi$ regardless of the decay exponent; it is the two sum rules $\sum a_{k}k=\sum a_{k}k^{2}=0$ that then isolate this quantity as the sole contributor to $g(z)$ at leading order. DMRG calculations on the $t$-$V$ model at four system sizes (L=64L=64512512, w=2w=244, $K\in[0.75,1.50]$) show that the ratio $R(K)=I_{3}(\Delta)/I_{3}(0)$ converges toward unity at small $z$: |1R|=0.015|1-R|=0.015 at $z=0.39$, $L=512$. This is consistent with $c_{K}=c$ for all $K$, though a residual $K$-dependence at the percent level cannot be excluded from the present data. At larger $z$, the convergence is slower ($|1-R|\approx 0.07$ at $z=0.79$, $L=512$), and the full function $g_{K}(z)$ may acquire $K$-dependent corrections through the subleading Slepian eigenvalues. A definitive determination requires infinite-system DMRG or periodic boundary conditions to eliminate open-boundary artifacts.
Open question. The product $R_{n}$ in Eq. (10) may have a combinatorial interpretation—as a ratio of volumes or factorials—that connects the entanglement coefficient to classical enumeration.
9 Conclusions
We have established a systematic analytical framework for the tripartite information of free fermions on two-dimensional lattices. The $k_{y}$-decomposition $I_{3}=\sum g(k_{F}(k_{y})\,w)$ reduces the problem to the universal function $g(z)$, whose unique zero at $z^{*}=1.329$ controls scale-dependent MMI. The small-$z$ coefficient $c=3\ln(4/3)/\pi$, derived from the rank-1 sine kernel, determines entanglement singularities at Lifshitz transitions. The $n$-partite generalization $c_{n}=(n/\pi)\ln R_{n}$ and the Rényi analysis—showing that von Neumann entropy uniquely gives linear Lifshitz sensitivity—complete the analytical picture.
Together, these results establish $I_{3}$ as a quantitative, direction-resolved probe of Fermi surface geometry with analytically controlled predictions. DMRG calculations on the $t$-$V$ model provide evidence that $c$ is robust against interactions.
Note.—Python code reproducing all numerical results is included as an ancillary file with the arXiv submission.
Appendix A Properties of $g(z)$: proof of the Proposition
(a) At $z=0$, all eigenvalues vanish, $h(0)=0$, so $g(0)=0$. For $z>0$: the rank-1 derivation (Section 3) gives $g(z)/z\to c$ as $z\to 0^{+}$. This is analytical.
(b) Numerically verified by evaluating $g^{\prime}(z)$ via central differences at 200 equally spaced points on $[0,0.60]$ using Toeplitz matrices of size $3w$ with $w=64$: $g^{\prime}(z)>0$ on $(0,0.56)$ and $g(0.56)\approx 0.088$.
(c) Numerically verified that $g^{\prime}(z)<0$ at 200 points on $[0.56,1.40]$ with the same method. Combined with $g(0.56)>0$ and $g(1.40)<0$, the intermediate value theorem gives a unique zero $z^{*}\in(0.56,1.40)$. The zero is located by bisection: $z^{*}(w{=}32)=1.32829$, $z^{*}(w{=}64)=1.32868$, $z^{*}(w{=}128)=1.32878$. The quoted value $z^{*}=1.329\pm 0.001$ is a conservative estimate: the variation between $w=64$ and $w=128$ is $0.0001$, but the extrapolation to $w\to\infty$ introduces additional uncertainty of order $1/w\sim 0.01$, which we round conservatively.
(d) We verify $g(z)<0$ on $(z^{*},10)$ by evaluating $g$ on grids of 500, 200, and 250 points on $z^{*},1.40]$, $[1.40,3.0]$, $[3.0,10.0]$ respectively, with $w=64$. The maximum of $g$ on any subgrid is bounded by the grid value plus the grid spacing times a numerically estimated derivative bound. For $z>10$, the asymptotic $g(z)\to(1/6)\ln(3/4)<0$ follows analytically from the Widom formula: the linear terms cancel by $\sum a_{k}k=0$ and the surviving logarithmic terms give $g_{\log}=(1/6)\ln(3/4)<0$.
Appendix B Neck transition derivation
Starting from $k_{F}(q)=\gamma\sqrt{q^{2}+\delta/\beta}$ and differentiating:
$\frac{\partial I_{3}}{\partial\delta}=\frac{Lw}{2\pi}\int_{-\Lambda}^{\Lambda}g^{\prime}(k_{F}w)\,\frac{\partial k_{F}}{\partial\delta}\,dq\approx\frac{cLw\gamma}{4\pi\beta}\cdot 2\,\mathrm{arcsinh}\!\left(\Lambda\sqrt{\beta/\delta}\right).$ (16)
For $\delta\to 0^{+}$: $\mathrm{arcsinh}(\Lambda\sqrt{\beta/\delta})\approx-\frac{1}{2}\ln\delta+\mathrm{const}$, confirming Eq. (15).